A bus was travelling with 48 passengers. When it arrived at the 1st stop, some passengers got off the bus and 3 people got on the bus. At the next stop half the passengers got off the bus and 7 people got on the bus. There were now 22 passengers on the bus.
1. Write down the equation from the statement
2. Find the number of passengers who came out of the bus as the 1st stop.

Answer by ptaylor(1894)


Let x=number of passengers that got off at first stop
Passengers left on board after first stop=(48-x)+3=45-x
At the next stop (1/2)(45-x) got off and seven got on, so:
Passengers left on board after second stop=(1/2)(45-x)+7
Now we are told that the passengers left on board after the second stop=22,so:
(1/2)(45-x)+7=22 multiply each side by 2
45-x+14=44 simplify
59-x=44 subtract 59 from both sides
59-59-x=44-59 collect like terms
-x=-15 divide both sides by -1
x=15--------------number of passengers that got off at first stop

CK
(1/2)(45-x)+7=22
(1/2)(45-15)+7=22
15+7=22
22=22

Hope this helps---ptaylor